OpenStudy (anonymous):

Which ordered pair is a solution of the equation y=-7x+2? (1,2) (8,-54) (5,-35) (1,-7)

OpenStudy (anonymous):

(8, -54)

terenzreignz (terenzreignz):

Just test it out. Replace x with the left-coordinate, and replace y with the right-coordinate. If all checks out, then that's the answer. Say I want to try (2, -12) Replace x with 2 and y with -12 -12 = -7(2) + 2 -12 = -14 + 2 -12 = -12 All checks out, so (2,-12) is a solution.